The four tires of an automobile are inflated to an absolute pressure of 2.00 x 105 Pa. Each tire has an area of 0.0240 m2 in contact with the ground. Determine the weight and mass of the automobile
1
Expert's answer
2012-03-06T11:51:00-0500
As it follows from the third Newton's law, the gravity force is equal in magnitude to the pressure force that acts on the tires. The pressure force is
F = P*S = 2.00 x 10^5 * 0.0240 = 4800 N.
So, the weight of an automobile is 4800N and it's mass (according to the second Newton's law) is M = 4800/g = 4800/9.8 ≈ 489.7959 kg.